UNDERSTANDING UO_{2}^{+} BY MODELING ELECTRONIC AND VIBRATIONAL SPECTROSCOPY

\underline{MICHAEL K. MROZIK}, RUSSELL M. PITZER, ANNE B. MCCOY, The Ohio State University, Columbus, OH 43210; BRUCE E. BURSTEN, University of Tennessee, Knoxville, TN 37996.

Spin-orbit configuration interaction (SO-CI) continues to be a valuable electronic structure tool in understanding the electronic states of small actinide containing molecules.^{[1]} Recent interest has a variation of a familiar actinide molecule, UO_{2}^{+}. The additional electron would appear to be a simple alteration or perturbation to an understood system, but experimentally it presents some questions for theory to investigate. Principal questions involve characterization of the lowest electronic states in UO_{2}^{+} involves occupation of the either the 5f_{ϕ} or the 5f_{δ} non-bonding orbitals and some simple excitation generates a multitude of electronic states. Use of self-consistent field with improved virtual orbital calculations (SCF) will again be the foundation of the SO-CISD determination of the potential energy surfaces (PES) of the ground and low lying excited states of the UO_{2}^{+} molecule. PES evaluation will occur along the internal coordinates of r_{UO} and θ_{O̸UO}, and will be variationally fit to determine bound vibrational levels present within the ground and low-lying excited states. Vibrational and electronic energy levels should help clarify and describe new and interesting spectra and other properties. [1]Gibson, J. K.; Haire, R. G.; Marcalo, J.; de Matos, A. P.; Mrozik, M. K.; Pitzer, R. M.; Bursten, B. E. Organomet. 2007, 26(16), 3947–3956.